what will the graph of |x| + [x] look like?

Grade:12

1 Answers

Nishant Vora IIT Patna
8 years ago
Dear student,

You can break the defination of the function in various intervals

|x| + [x] = \left\{\begin{matrix} x+ 0 & 0\leq x<1 & \\ x + 1 & 1\leq x<2 & & \end{matrix}\right. and so on
then plot in respective interval